# Evaluating Small Language Models for News Summarization: Implications and Factors Influencing Performance

### v1

@@ Abstract
Borui Xu Shandong University boruixu@mail.sdu.edu.cn                        Yao Chen National University of Singapore yaochen@nus.edu.sg                        Zeyi Wen HKUST (Guangzhou) HKUST wenzeyi@ust.hk
Weiguo Liu Shandong University weiguo.liu@sdu.edu.cn                        Bingsheng He National University of Singapore hebs@comp.nus.edu.sg
@@ 1 Introduction
First, How do different SLMs compare to each other, and how do they compare to LLMs? We benchmark 19 SLMs on 2,000 news samples. Our findings reveal that the news summarization capabilities of different SLMs vary significantly as shown in Figure 1. And the best-performing SLMs can generate news summaries comparable in quality to those produced by LLMs while producing shorter summaries. Notably, Phi3-Mini and Llama3.2-3B-Ins emerge as the top performers.
@@
The rest of the paper is organized as follows: Section 2 covers the background and related work, Section 3 introduces our reference summary generation, Section 4 illustrates the benchmark design, Section 5 presents the benchmarking results of SLMs, Section 6 analyzes the influencing factors on SLM summarization, and Section 7 concludes with a summary of findings and future work.
@@ 2.2 Text Summarization and Its Evaluation
Reference-based evaluation is a commonly used method. It scores generated text by calculating its similarity to reference texts using various metrics, and many such metrics have been proposed Lin (2004); Papineni et al. (2002); Sellam et al. (2020); Zhang et al. (2020b); Yuan et al. (2021). In addition to human evaluation, Huang et al. (2020) and  Fabbri et al. (2021) also used many reference-based metrics to evaluated models to and found metrics like Rouge and BertScore are very close to human evaluations. Maynez et al. (2023) used the Rouge metric to measure text similarity. Tiny Titans Fu et al. (2024) assessed the performance of small models in meeting summarization and found FLAN-T5 performed best.
@@
In conclusion, existing evaluations either focus on the news summarization capabilities of traditional models or LLMs, resulting in a gap in SLM text summarization.
@@ 5.5 Human Evaluation
To further validate the effectiveness of our metrics, we performed a small-scale human evaluation, referencing existing studies Fabbri et al. (2021). Specifically, six highly educated annotators evaluated the relevance of news summaries generated by five SLMs based on the original news articles (scored on a 1-5 scale, with higher scores indicating greater relevance). We randomly selected 20 news instances from the BBC2024 dataset as our evaluation set. As shown in Table 7, The results demonstrate a strong alignment between BERTScore and human evaluations, achieving a Kendall correlation coefficient of 1.

### v2

@@ Abstract
Borui Xu Shandong University boruixu@mail.sdu.edu.cn                        Yao Chenthanks: Corresponding authors National University of Singapore yaochen@nus.edu.sg                        Zeyi Wen HKUST(GZ) & HKUST wenzeyi@ust.hk
Weiguo Liu* Shandong University weiguo.liu@sdu.edu.cn                        Bingsheng He National University of Singapore hebs@comp.nus.edu.sg
@@ 1 Introduction
First, How do different SLMs compare to each other, and how do they compare to LLMs? We benchmark 19 SLMs on 2,000 news samples [^2] The codes and results are available at https://github.com/Xtra-Computing/SLM_Summary_Benchmark. Our findings reveal that the news summarization capabilities of different SLMs vary significantly as shown in Figure 1. And the best-performing SLMs can generate news summaries comparable in quality to those produced by LLMs while producing shorter summaries. Notably, Phi3-Mini and Llama3.2-3B-Ins emerge as the top performers.
@@
The rest of the paper is organized as follows: Section 2 covers the background and related work, Section 3 introduces reference summary generation, Section 4 illustrates the benchmark design, Section 5 presents presents the benchmarking results of SLMs, Section 6 analyzes the influencing factors on SLM summarization, and Section 7 concludes our findings.
@@ 2.2 Text Summarization and Its Evaluation
Reference-based evaluation is a commonly used method. It scores generated text by calculating its similarity to reference texts using various metrics, and many such metrics have been proposed Lin (2004); Papineni et al. (2002); Sellam et al. (2020); Zhang et al. (2020b); Yuan et al. (2021). In addition to human evaluation, Huang et al. (2020) and  Fabbri et al. (2021) also used many reference-based metrics to evaluated models and found metrics like Rouge and BertScore are very close to human evaluations. Maynez et al. (2023) used the Rouge metric to measure text similarity. Tiny Titans Fu et al. (2024) assessed the performance of small models in meeting summarization and found FLAN-T5 performed best.
@@
In conclusion, existing evaluations primarily evaluate the news summarization capabilities of traditional models or LLMs, leaving a gap in the evaluation of SLMs.
@@ 5.5 Human Evaluation
To further validate the effectiveness of our metrics, we performed a small-scale human evaluation, referencing existing studies Fabbri et al. (2021). Specifically, six highly educated annotators evaluated the relevance of news summaries generated by five SLMs based on the original news articles (scored on a 1-5 scale, with higher scores indicating greater relevance). We randomly selected 20 news instances from the BBC2024 dataset as our evaluation set. As shown in Table 7, the average scores of each model demonstrate a strong alignment between BertScore and human evaluations, achieving a Kendall correlation coefficient of 1.
@@ 7 Conclusion
Acknowledgments
This work is funded by the National Key R&D Program of China (No. 2019YFA0709400), the NSFC Project (No. 62306256), the Guangzhou Science and Technology Development Projects (No. 2023A03J0143 and No. 2024A04J4458), and the Guangzhou Municipality Big Data Intelligence Key Lab (NO. 2023A03J0012). This work is also supported by the National Research Foundation, Singapore under its Industry Alignment Fund – Pre-positioning (IAF-PP) Funding Initiative. Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of National Research Foundation, Singapore. Borui’s work was done when he was a visiting student at the National University of Singapore.